OpenStudy (anonymous):

What mass in grams of CO2 is produced if 35.0 moles of magnesium carbonate are heated

OpenStudy (anonymous):

first write the chemical reaction equation, then from equation see the ratio between MgCO3 and CO2 and from there you need to take ratio in mind and put your moles in n=m/M this equation where m(CO2)= n*M. your answer is: 1540 g